You would like to start a business and have decided to move back into your parents’ basement as you save up the $50,000 you will need in starter costs. You procure a job that will allow you to contribute $550 into a savings account semi monthly (24 times a year). The account will be compounded at that same interval and will earn an annual interest rate of 3.259%. The instrument only allows you to collect at the end of complete years. How many years will you need to invest your money to have over your goal and how much will you have?
Answer #1

FROM THIS QUESTION IT IS UNDERSTOOD THAT WE HAVE TO FIND OUT THE COMPOUND INTEREST OF $550 AT THE RATE OF 3.259% ON ANNUAL BASIS.

YOU NEED 3 AND HALF YEARS INVESTMENT TO REACH THE DESIRED GOAL

IF YOU INVEST $1100 A MONTH AND IT REACH $13200 A YEAR

APART FROM $13200 AN INTEREST RATE COMPOUNDED ON SEMI MONTHLY BASIS WILL ALSO RECEIVE FOR THIS AMOUNT

WHICH MEANS THAT FIST DEPOSIT OF $550 WILL GET INTEREST AT THE RATE OF 3.259% ON SEMI MONTHLY BASIS ON 15 TH OF THE MONTH (ASSUMING THAT THE AMOUNT DEPOSITED ON 1 ST DAY OF THE MONTH)

($550*3.259/100)*0.5/12=0.74          

HERE 0.5/12 MEANS SEMI MONTHLY INTEREST.

TO FIND OUT THE NUMBER OUT HOW MANY YEARS NEEDED TO GET THE DESIRED AMOUNT YOU CAN USE THE COMPOUND INTEREST FORMULA

A=P (1 + r/n) (nt)

WHERE P IS THE PRINCIPAL AMOUNT ,IE $550

R IS THE RATE OF INTEREST =3.259/100

N IS THE NUMBER OF TIME DEPOSIT = 24

T IS THE TIME PERIOD OF DEPOSIT WHICH IS NOT MENTIONED IN THIS QUESTION

SO USE THIS INFORMATION AND CALCULATE THE ANSWER YOURSELF .TRY IT .ALL THE BEST

THANK YOU